Please help me solve this :
If an object is thrown vertically upward from the ground with initial speed of 160 feet per second, its height (in feet) t seconds later is given by the function H(t) = -16t^2 + 160t
a) graph the function H(t)
b) use the graph in part a) to determine when the object will hit the ground
c) use the graph in part (a) to estimate how high the object will rise.
I think b) 4 seconds
c) 5 feet
Thank you for sending me correct answer

Answer by scott8148(6628)

the graph is a downward opening parabola
the maximum height is at the vertex , on the axis of symmetry

the general equation for the axis of symmetry is ___ x = -b / (2a)
___ in this case ___ x = -160 / [2(-16)] = 5

H(5) = -16(5^2) + 160(5) = 400

so the max height is 400 ft and it takes 5 sec to get there

time up equals time down , so the "flight" time is 10 sec